OpenStudy (anonymous):

Which statement is true about the product square root of 2(3square root of 2 + square root of 14)?

OpenStudy (anonymous):

\[\sqrt{2}(\sqrt[3]{2}+\sqrt{18})\]

OpenStudy (anonymous):

the a is 3 sorry about that.

OpenStudy (anonymous):

my 4 options are It is rational and equal to 12. It is rational and equal to 18. It is irrational and equal to 6 + 2square root of 18. It is irrational and equal to 3 +square root of 18.

OpenStudy (kira_yamato):

√18 = √(2* 3^2) = 3√2 So the question is actually √2 * (3√2 + 3√2) = √2 * 6√2 = 12

OpenStudy (anonymous):

one secound im trying to proccess that in my head (hears static in background)

OpenStudy (anonymous):

K i got it wow didnt look at it that way you saved ma life thanks!